### v1

# UniVoice: Unifying Autoregressive ASR and Flow-Matching based TTS with Large Language Models

Abstract
Large language models (LLMs) have demonstrated promising performance in both automatic speech recognition (ASR) and text-to-speech (TTS) systems, gradually becoming the mainstream approach. However, most current approaches address these tasks separately rather than through a unified framework. This work aims to integrate these two tasks into one unified model. Although discrete speech tokenization enables joint modeling, its inherent information loss limits performance in both recognition and generation. In this work, we present UniVoice, a unified LLM framework through continuous representations that seamlessly integrates speech recognition and synthesis within a single model. Our approach combines the strengths of autoregressive modeling for speech recognition with flow matching for high-quality generation. To mitigate the inherent divergence between autoregressive and flow-matching models, we further design a dual attention mechanism, which switches between a causal mask for recognition and a bidirectional attention mask for synthesis. Furthermore, the proposed text-prefix-conditioned speech infilling method enables high-fidelity zero-shot voice cloning. Experimental results demonstrate that our method can achieve or exceed current single-task modeling methods in both ASR and zero-shot TTS tasks. This work explores new possibilities for end-to-end speech understanding and generation.
Introduction
Recent years have witnessed remarkable progress in large language models (LLMs), with systems such as GPT-4 (Achiam et al. 2023), LLaMA (Touvron et al. 2023), and Qwen (Bai et al. 2023) demonstrating unprecedented capabilities in understanding, generation, and reasoning. Modern LLMs have achieved broad knowledge representation and strong generalization in diverse domains.
This success has naturally been extended to speech processing, where researchers have adapted LLM frameworks to handle continuous speech signals through various discretization strategies (Chen et al. 2024a; Wang et al. 2024a; Du et al. 2023; Ma et al. 2024). The prevailing approach involves converting raw waveforms into discrete tokens using self-supervised learning representations (Hsu et al. 2021) or quantization based on neural codecs (Zeghidour et al. 2021). These tokenized representations enable the direct application of existing LLM architectures to speech tasks: for ASR tasks, Codec-ASR (Dhawan et al. 2024) performs comprehensive analysis on building ASR systems with discrete codes; for TTS tasks, systems such as VALL-E (Wang et al. 2023) and AudioLM (Borsos et al. 2023) formulate speech generation as discrete token sequence modeling. Concurrently, diffusion models (Ho, Jain, and Abbeel 2020) and flow matching models (Lipman et al. 2022; Liu, Gong, and Liu 2022) have rapidly gained prominence in speech synthesis, establishing new benchmarks for generation quality. Pioneering works like Grad-TTS (Popov et al. 2021) and F5-TTS (Chen et al. 2024c) demonstrate their exceptional ability to directly model continuous speech representations, achieving unprecedented fidelity that has made them increasingly prevalent in recent research.
Several pioneering efforts have attempted to unify ASR and TTS within single LLM frameworks using discrete representations (Wang et al. 2024a; Du et al. 2023; Tian et al. 2025). Viola (Wang et al. 2024a) converts speech utterances to discrete tokens using an offline neural codec encoder and treats all tasks as token-based sequence prediction problems, while LauraGPT (Du et al. 2023) employs a novel data representation combining continuous and discrete features for audio signals. However, such methods face inherent limitations; quantization unavoidably eliminates perceptually critical acoustic details, driving our investigation of continuous representation alternatives.
We present UniVoice, a novel architecture unifying ASR and TTS within a continuous signal space while retaining LLM framework scalability. The architecture maintains continuous representations across both tasks, employing autoregressive modeling (AR) for ASR to leverage its sequential prediction strengths, while utilizing flow matching (FM) for TTS to capitalize on its high-fidelity generation advantages. To resolve the inherent incompatibility between AR’s causal masking and FM’s non-autoregressive requirements, we design a dual attention mask mechanism that switches between causal masking for recognition and bidirectional attention for synthesis. Furthermore, our text-prefix guided speech infilling method enables high-fidelity zero-shot voice cloning.
We evaluated the ASR and TTS performance of our method on the LibriHeavy dataset (Kang et al. 2024). Experiments demonstrate that our approach achieves speech synthesis performance comparable to that of current state-of-the-art methods while maintaining competitive speech recognition capability.
In summary, the main contributions of this work are as follows.
  • We present a unified framework that integrates autoregressive speech recognition with flow-matching-based synthesis in pre-trained LLMs, enabling joint speech understanding and generation within a single model.
  • We propose a dual-attention mechanism that switches between causal masking for recognition and bidirectional attention for synthesis, along with text-prefix guided infilling for high-fidelity zero-shot voice cloning.
  • Extensive experiments demonstrate that our unified approach matches or surpasses state-of-the-art specialized models in both speech recognition and zero-shot speech synthesis performance, while maintaining parameter efficiency.
Related Work
Speech Language Models
The emergence of ChatGPT has demonstrated the remarkable capabilities of LLMs, inspiring significant advancements in audio and speech processing. Recent research has successfully adapted LLMs to model discrete speech tokens, enabling high-quality speech generation and editing. Several notable approaches have emerged in this field (Wang et al. 2023; Anastassiou et al. 2024; Meng et al. 2024; Ye et al. 2025). Wang et al. (2023) introduced VALL-E, framing TTS as a conditional audio codec language modeling task with in-context learning capabilities. This work was subsequently enhanced by Chen et al. (2024a) through repetition-aware sampling and grouped codec modeling in VALL-E 2. Copet et al. (2024) developed MusicGen, utilizing delay patterns to model multiple parallel streams of audio tokens. Similarly, Peng et al. (2024) presented VoiceCraft, which combines delay patterns with causal masking for zero-shot speech editing and synthesis. Jiang et al. (2023) proposed Mega-TTS, a zero-shot TTS system that decomposes mel-spectrograms into content, timbre, prosody, and phase attributes, modeling each component according to its intrinsic properties. Yang et al. (2023) developed UniAudio, an audio foundation model capable of handling multiple generation tasks through multi-scale transformer-based codec modeling (Yu et al. 2023). Seed-TTS (Anastassiou et al. 2024) proposes utilizing both speech tokenizer based language model and diffusion acoustic modeling the first time for natural and high-quality speech generation. Similarly, CosyVoice (Du et al. 2024) proposes supervised semantic tokens for audio codec modeling and flow matching for acoustic details modeling. MaskGCT (Wang et al. 2024b) designs a masked generative codec transformer for zero-shot speech synthesis.
In addition to using LLM for modeling speech synthesis, there are also some works applying LLM for speech recognition or understanding tasks (Bai et al. 2024; Ma et al. 2024). SALMONN  (Tang et al. 2023) uses a dual encoder from Whisper speech encoder (Radford et al. 2023) and BEATS audio encoder (Chen et al. 2022b) for speech and audio understanding. SeedASR (Bai et al. 2024) is developed under the audio conditional LLM framework, utilizing the capabilities of LLM by submitting continuous speech representations, instructions, and contextual information into LLM.
Diffusion based Speech Generative Model
The diffusion (Ho, Jain, and Abbeel 2020; Song et al. 2020) models first achieved great success in the field of image generation, and then there are many works in the field of speech and audio using diffusion or flow matching (Lipman et al. 2022) to model speech and audio generation. Due to the advantages of diffusion and flow matching in modeling continuous features, diffusion- or flow-based audio generative models can often generate high-quality audio (Liu et al. 2022; Guan et al. 2024a; Anastassiou et al. 2024; Jiang et al. 2025; Wang et al. 2025b, a). Voicebox (Le et al. 2024) proposes a text-guided speech infilling task to generate masked speech segments based on surrounding audio and a provided text transcript. Matcha-TTS (Mehta et al. 2024) and VoiceFlow (Guo et al. 2024b) utilize the optimal transport conditional flow matching based on Grad-TTS (Popov et al. 2021) for high-quality and faster TTS. ReFlow-TTS (Guan et al. 2024b) models the Ordinary Differential Equation (ODE) based on Rectified Flow (Liu, Gong, and Liu 2022) for high-fidelity and efficient speech synthesis. NaturalSpeech2 (Shen et al. 2023) models zero shot speech synthesis as a conditional latent diffusion model with codec-latent embeddings and attention-based in-context modeling. FlashSpeech (Ye et al. 2024) develops a large-scale zero-shot speech synthesis with a latent consistency model and adversarial consistency training approach for efficient zero-shot speech synthesis. F5-TTS (Chen et al. 2024c) is a fully non-autoregressive text-to-speech system based on flow matching with Diffusion Transformer for zero-shot TTS. Some works also use latent diffusion or a latent flow matching model for text-to-audio generation (Liu et al. 2023; Ghosal et al. 2023; Guan et al. 2024c).
Figure 1: An overview of UniVoice model. Blue elements (blocks and lines) denote ASR components, while green elements represent TTS components. The gradient-colored modules (blue-to-green or green-to-blue) indicate shared components between both ASR and TTS systems.
Unified Models for Speech Processing
With the development of language models, they can handle various downstream tasks as a unified model in the field of text processing (Raffel et al. 2020; Brown et al. 2020). Some works have begun to explore the use of language models as the backbone of the unified model for processing speech and text tasks (Ao et al. 2021; Rubenstein et al. 2023; Zhang et al. 2023b), thus completing various related tasks in the speech modality. SpeechNet (Chen et al. 2021) and SpeechT5 (Ao et al. 2021) perform various speech tasks with an encoder-decoder model, specifically SpeechT5 needs to pre-train first and then fine-tune on subsequent tasks. Viola (Wang et al. 2024a) follows the VALL-E paradigm and integrates speech recognition, machine translation, and speech synthesis into a unified codec language model. LauraGPT (Du et al. 2023) encodes the input audio into continuous representations using an audio encoder and generates the output audio from discrete codec codes. OpusLM (Tian et al. 2025) is designed to accept and generate multistream discrete tokens in both text and speech modalities with pre-trained LLM. We propose UniVoice, which uses continuous speech representations as input and output features and combines autoregression and flow matching in one transformer, to make it more suitable for speech recognition and speech synthesis, respectively.
Note that the primary focus of this work is developing a unified speech processing model, and we do not explore speech dialogue applications (Zhang et al. 2023a; Xie and Wu 2024; Fu et al. 2024; Chen et al. 2024b) that require joint speech understanding and generation. Therefore, such tasks are beyond the scope of this paper.
Preliminaries
Autoregressive Language Modeling
Autoregressive language modeling is a fundamental task in natural language processing that aims to predict the next word or character in a sequence based on the preceding context. This approach assumes that the probability distribution of the current word depends solely on the preceding words.
The core objective of autoregressive language modeling is to estimate the conditional probability of the next word in a sequence, given the preceding words. This can be mathematically represented as:
$P(x_{1},x_{2},...,x_{n})=\prod_{i=1}^{n}p_{\theta}(x_{i}|x_{<i})$ (1)
where $x=x_{1},x_{2},...,x_{n}$ represents the sequence of words. This formulation describes an autoregressive language modeling task, in which the model aims to predict the probability distribution of each token $x_{i}$ based on the preceding tokens $x_{<i}$ in the sequence. This prediction is made using a probability distribution $p_{\theta}$ that is parameterized by $\theta$. The model is trained by minimizing the cross-entropy loss between the predicted distribution $p_{\theta}$ and the empirical distribution of the training data. This optimization process results in the language model loss, which can be expressed as:
$L_{LM}(\theta)=E_{x_{i}}[-\log p_{\theta}(x_{i}|x_{<i})]$ (2)
Flow Matching
The objective of Flow Matching (FM)(Lipman et al. 2022) is to construct a flow that transforms a sample $X_{0}\sim p$, drawn from a source distribution $p$, into a target sample $X_{1}=\psi_{1}(X_{0})$ such that $X_{1}\sim q$ conforms to a desired distribution $q$.
The goal of FM is to learn the parameters $\theta$ of a velocity field $u_{t}^{\theta}$, which is implemented using a neural network. We define the source distribution as $p_{0}=\mathcal{N}(x|0,1)$, a standard normal distribution. The probability path $p_{t}$ is then constructed as the aggregation of conditional probability paths $p_{t|1}(x|x_{1})$, each of which is conditioned on one of the data examples $X_{1}=x_{1}$ from the training dataset. This path is also known as the conditional optimal transport path. Using this probability path, we may define the random variable $X_{t}\sim p_{t}$ by drawing $X_{0}\sim p$ and $X_{1}\sim q$, and taking the linear combination $X_{t}=tX_{1}+(1-t)X_{0}\sim p_{t}$.
We aim to regress our velocity field $u_{t}^{\theta}$ towards a target velocity field $u_{t}$ that is known to generate the desired probability path $p_{t}$. To achieve this, the Flow Matching loss is defined as:
[…]
The objective mentioned above is challenging to implement in practice because $u_{t}$ is a complex function that governs the joint transformation between two distributions. Fortunately, this objective can be significantly simplified by conditioning the loss on a single target example $X_{1}=x_{1}$ randomly selected from the training dataset. And then we can establish a feasible FM loss, known as the conditional FM loss:
$L_{CFM}(\theta)=E_{t,X_{t},X_{1}}||u_{t}^{\theta}(X_{t})-u_{t}(X_{t}|X_{1})||^{2}$ (4)
Finally, using the conditional optimal transport path, we can derive the OT-CFM loss:
[…]
UniVoice
Model
As depicted in Figure 1, our framework features a dual-branch hybrid architecture comprising: (1) a causal transformer for ASR tasks, and (2) a flow-matching-based Diffusion Transformer for TTS synthesis.
ASR
The ASR component processes input speech through an audio encoder module, which consists of: 1) a Whisper encoder (Radford et al. 2023) for the extraction of speech features and 2) an adapter network for semantic alignment. The inherent structure of text as a linear sequence aligns well with the standard 1D positional embeddings of the LLM, which are sufficient for text modeling and speech understanding tasks.
TTS
The TTS component accepts three distinct inputs: 1) transcript token sequence (text condition), 2) noisy speech features, and 3) masked speech features. These inputs are processed through the following pipeline: 1) the noisy and masked speech features are concatenated along the feature dimension, 2) the transcript token sequence is prepended to the concatenated speech features, 3) the flow step $t$, embedded via sinusoidal positional encoding, is inserted between the text and speech sequences as a conditioning signal.
Compared to F5-TTS (Chen et al. 2024c), we replace the adaLN-zero with the in-context modeling approach of DiT to maintain the original LLM structure. Additionally, we propose a strategic sequence organization (text prefix) to preserve the in-context learning abilities inherent in LLMs.
As illustrated in Figure 2, we introduce two variants of the TTS model for systematic comparison: (1) our primary UniVoice-TTS-infilling model that implements the voice cloning task of our UniVoice framework and (2) UniVoice-TTS-speaker, a simplified architecture designed for baseline evaluation. For UniVoice-TTS-speaker, we adopt a direct mel-spectrogram generation approach, bypassing masked infilling paradigms entirely. The system produces mel-spectrograms conditioned on input text embeddings while utilizing speaker embeddings to control vocal characteristics, enabling effective voice cloning. In this setting, for the speaker encoder, we utilize the first layer of the XLSR-53 model (Conneau et al. 2020) to extract a global embedding that effectively captures the speaker’s timbre characteristics.
Figure 2: Two variants of TTS model designs in UniVoice. (a) uniVoice-TTS-speaker. (b) UniVoice-TTS-infilling.
Training
ASR
For automatic speech recognition, we employ an autoregressive objective function to optimize the model. The architecture follows a standard transformer-based language model, with the key modification being the incorporation of an audio encoder and an adapter. The transformer processes the audio encoder embeddings as input sequences, while the output embeddings are projected through an autoregressive prediction head to generate token probabilities.
TTS
We formulate speech generation as a text-prefix-guided speech-infilling task, where the model predicts speech segments conditioned on both surrounding audio context and transcript text provided as prefix. Within the conditional flow matching framework, the model receives two key inputs: noisy speech representation: $tX_{1}+(1-t)X_{0}$ and masked speech features: $(1-m)\odot X_{1}$, where $m$ denotes a binary mask with identical dimensionality to $X_{1}$.
Text transcripts are tokenized using a pre-trained language model tokenizer to produce discrete tokens $z$, which are prefixed to the speech input sequence. The model learns to reconstruct the masked portions $m\odot X_{1}$ conditioned on both the unmasked context $X_{ctx}=(1-m)\odot X_{1}$ and text tokens $z$. This corresponds to approximating the target distribution $p_{1}$ through the conditional probability $P(m\odot X_{1}|(1-m)\odot X_{1},z)$, which converges to the true data distribution $q$ during training. Finally, the OT-CFM loss in the scenario is:
[…]
Unified Training Objective
The complete loss function combines the ASR and TTS objectives through a weighted sum.
$L_{total}=\lambda L_{LM}(\theta)+L_{audio}^{cfm}(\theta)$ (7)
where $L_{LM}(\theta)$ represents the autoregressive loss for ASR, $L_{audio}^{cfm}(\theta)$ denotes the optimal transport-based conditional flow matching loss for TTS, $\lambda$ serves as a balancing hyperparameter between the two objectives.
Following F5-TTS (Chen et al. 2024c), we also apply Classifier-Free Guidance (CFG) by dropping the condition at a certain rate during training.
Inference
ASR
The ASR inference follows a conventional autoregressive decoding process. The trained transformer model sequentially predicts token probabilities, generating the output sequence through iterative sampling of the most probable next token conditioned on the preceding tokens.
TTS
For TTS generation, the model requires three key inputs: 1) reference audio mel-spectrogram features ($X_{ref}$) providing paralinguistic information, 2) reference transcription ($Y_{ref}$) establishing baseline duration patterns and providing linguistic content of the reference audio, and 3) target text prompt ($Y_{gen}$) providing the target linguistic content.
Following the F5-TTS approach (Chen et al. 2024c), we compute the duration ratio as:
$duration\_ratio=\frac{len(Y_{gen})}{len(Y_{ref})}$ (8)
The generation process involves: 1) tokenizing $Y_{ref}$ and $Y_{gen}$ to form the text condition $z$, 2) using $X_{ref}$ as the acoustic context $X_{ctx}$, 3) initializing from noise $X_{0}$ and solving the flow ODE to obtain $X_{1}$.
The ODE solver performs numerical integration along the learned flow path, with step sizes determined by the Number of Function Evaluations (NFE) parameter. After generating the mel-spectrogram, the reference portion $X_{ref}$ is discarded, and the synthesized mel-spectrogram is converted to waveform using a neural vocoder.
Attention Mask Design
When training the two tasks simultaneously, we need to pay special attention to the setting of attention masks for different tasks. UniVoice needs to meet the requirements for simultaneously training speech understanding and generation tasks. As illustrated in Figure 1, for ASR tasks, we use the same causal mask as the original LLMs.
For TTS tasks, we use a bidirectional attention mask and utilize the characteristics of LLM text modeling to model the text content in TTS tasks, similar to in-context learning in VALL-E (Wang et al. 2023).
Table 1: Performance Comparisons of UniVoice and prior works. A-WER-clean, A-WER-other represents the ASR WER evaluation result on LibriSpeech test-clean dataset and LibriSpeech test-other dataset, respectively. represents that the results are reported in NaturalSpeech3(Ju et al. 2024).
[…]
Ground Truth - - 0.69 2.43 4.07 +0.09 3.82 - -
Unified Models SpeechT5 0.140.14B 0.96K 0.33 5.91 3.32 -0.28 3.35 4.4 10.4
LauraGPT 2.02.0B 60K - 8.62 - - - 4.4 7.7
OpusLM-0.4B 0.40.4B 213K - 19.8 - - - 4.2 8.7
OpusLM-7B 77B 213K - 4.60 - - - 2.3 5.2
UniVoice (Ours) 0.4B 50K 0.56 4.06 3.72 0.00 3.88 3.0 6.3
Only Zero-shot TTS Models CosyVoice 0.40.4B 170K 0.66 3.59 4.17 +0.06 3.96 - -
MaskGCT 1.01.0B 100K 0.66 2.49 3.85 +0.04 3.92 - -
F5-TTS 0.30.3B 100K 0.66 2.54 3.84 +0.03 3.90 - -
FireRedTTS 0.60.6B 248K 0.47 2.69 3.91 -0.01 3.85 - -
VALL-E 0.30.3B 60K 0.47 6.11 3.68 - - - -
NaturalSpeech2 0.40.4B 60K 0.55 1.94 3.88 - - - -
UniVoice-TTS (Ours) 0.40.4B 50K 0.56 4.66 3.92 +0.02 3.86 - -
Only ASR Models Whisper-small 0.20.2B 680K - - - - - 3.4 7.6
Whisper-large-v2 1.51.5B 680K - - - - - 2.7 5.2
Whisper-large-v3 1.51.5B 680K - - - - - 1.9 3.6
Whisper-large-v3-turbo 0.80.8B 680K - - - - - 1.9 3.5
Paraformer 0.20.2B 20K - - - - - 3.5 8.2
Zipformer 0.150.15B 0.96K - - - - - 2.0 4.4
UniVoice-ASR (Ours) 0.40.4B 50K - - - - - 2.5 4.2
Experiments
Experimental Setups
Dataset
We conduct the training of the proposed UniVoice  using the LibriHeavy (Kang et al. 2024) dataset for both ASR and TTS training, which contains a total of about 50K hours of duration, which are sampled at 16kHz. In practice, we upsample the sample rate to 22.05kHz. For the zero-shot TTS evaluation, we use the LibriSpeech-PC test set the same as in F5-TTS (Chen et al. 2024c). For the ASR task, we directly use the LibriSpeech test-clean and test-other subsets for evaluation. We extract the 80-bin mel-spectrogram with a frame size of 1024 and a hop size of 256.
Implementation details
We use SmolLM2-360M (Allal et al. 2025) as the underlying language model for initialization. We use the Whisper-large-v3-turbo encoder as the audio encoder, and we utilize an adaptive average pooling layer as the adapter for downsampling. For the vocoder, we use BigvGAN (Lee et al. 2022) in practice.
UniVoice is trained for 10 epochs using the AdamW optimizer with a learning rate of 1.5e-3 and the cosine scheduler, $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.95, the warmup step is 20000 steps. All models are trained with a batch size of 160000 audio frames in total. Considering using LLM as the backbone, the ASR task is simpler compared to the TTS task, the $\lambda$ in the loss function is set to a small value 0.005. A random 70% to 100% of mel frames is masked in the training of the TTS model. For CFG training, we randomly omit text tokens with a drop probability of 0.2, and masked speech is dropped with a probability of 0.3. For TTS inference, the value of the CFG weight is set to 2, and the inference step is set to 32.
Evaluation metrics.
For zero-shot TTS, we conduct a comprehensive evaluation, encompassing both objective and subjective measures, to assess the sample quality (UTMOS, CMOS), speaker similarity (SIM, SMOS) and robustness (WER). In specific, 1) for speech quality, we employ UTMOS [^1] https://github.com/sarulab-speech/UTMOS22 (Saeki et al. 2022), which is a surrogate objective metric of MOS; and we employ comparative mean option score (CMOS) to evaluate the sample naturalness subjectively; 2) for speaker similarity, we use a WavLM-large-based speaker verification model (Chen et al. 2022a) to extract speaker embeddings for calculating the cosine similarity of synthesized and ground truth speeches as SIM, and we employ similarity mean option score (SMOS) to evaluate the similarity subjectively; 3) for Word Error Rate (WER), we use an ASR model to transcribe generated speech. We use Whisper-large-v3 to compute the WER.
For ASR, we compare the WER of different systems evaluated on the LibriSpeech test-clean set and the test-other set using the Whisper-large-v3 ASR model.
Baselines
For unified models, we compare our UniVoice  with baselines: SpeechT5 (Ao et al. 2021), LauraGPT (Du et al. 2023), OpusLM (Tian et al. 2025).
For the zero-shot TTS task, we compare our UniVoice with baselines: VALL-E (Wang et al. 2023), NaturalSpeech2 (Shen et al. 2023), F5-TTS (Chen et al. 2024c), CosyVoice (Du et al. 2024), FireRedTTS (Guo et al. 2024a), MaskGCT (Wang et al. 2024b), UniVoice-TTS. UniVoice-TTS is the model trained only for TTS using our model framework.
For the ASR task, we compare our UniVoice  with baselines: Whisper (Radford et al. 2023) series, Paraformer (Gao et al. 2022), Zipformer (Yao et al. 2023), UniVoice-ASR. UniVoice-ASR is the model trained only for ASR using our model framework.
Details of compared models are in Appendix A.
Main Results
Experimental Results on Zero-shot TTS
We comprehensively evaluate our zero-shot TTS system in three critical dimensions: robustness, generation similarity, and generation quality.
Robustness Evaluation
We evaluated model robustness through WER measurements on LibriSpeech test-clean. As shown in Table 1, UniVoice demonstrates significant improvements over existing unified approaches, achieving a 12% relative WER reduction compared to the best unified baseline. This performance gain highlights the effectiveness of our dual attention mechanism and text-prefix-guided speech-infilling approach in maintaining speech intelligibility.
However, our analysis reveals an important trade-off: while outperforming other unified models, UniVoice shows a slight but consistent WER gap compared to specialized single-task TTS systems. We attribute this difference to two key factors inherent in unified architectures: 1) structural constraints imposed by shared parameters, which limit task-specific optimization. 2) the competing objectives of maintaining both recognition accuracy and generation quality.
Notably, UniVoice demonstrates a 13% WER drop over UniVoice-TTS, confirming that joint ASR-TTS training enhances speech intelligibility through shared linguistic representations.
Generation Similarity Analysis
We evaluate speaker similarity through both objective SIM scores and subjective SMOS tests (8 listeners, 20 utterances). Our results show that UniVoice outperforms FireRedTTS, NaturalSpeech2, and VALL-E, while showing a modest performance degradation compared to CosyVoice, MaskGCT, and F5-TTS. This performance gap likely stems from insufficient utilization of the AdaLN-zero modulation mechanism inherited from DiT models.
Interestingly, UniVoice achieves parity with UniVoice-TTS in similarity metrics, indicating that multitask learning preserves speaker characteristics effectively.
Generation Quality Assessment
Quality evaluation combines UTMOS scores and CMOS tests (8 listeners, 20 utterances). UniVoice sets a new state-of-the-art among unified models, with a 0.4 UTMOS improvement over previous approaches. However, we observe a 0.2 UTMOS degradation compared to UniVoice-TTS, suggesting that joint optimization involves subtle trade-offs in naturalness.
These results collectively demonstrate that while our unified architecture achieves remarkable performance across all metrics, there remains an inherent tension between multitask generalization and single-task optimization that warrants further investigation.
Experimental Results on ASR
We compare UniVoicewith the state-of-the-art neural network-based models, Whisper-small, Whisper-large-v2 and Whisper-large-v3, Whisper-large-v3-turbo, Paraformer and Zipformer. Moreover, we also compare UniVoicewith previous unified models. As shown in Table 1, the results indicate that UniVoice achieves an excellent level of audio comprehension, although it is a relatively small model trained on a relatively small dataset. Moreover, compared to UniVoice-ASR, UniVoice performs slightly worse due to the unified training of two distinct objectives.
In conclusion, UniVoice demonstrates a trade-off in performance: while it achieves improved robustness in TTS (as evidenced by lower TTS WER), it experiences a slight degradation in both TTS naturalness and ASR performance compared to corresponding single-task models.
Ablation Study of TTS model variants
We conduct a comprehensive comparison between two TTS variants within our UniVoice framework: (1) the proposed speech-infilling-based model and (2) a speaker-embedding-conditioned baseline. As detailed in Table 2, the speech-infilling-based approach demonstrates consistent superiority across all evaluation metrics, robustness (WER reduction of 18%), generation similarity (0.27 SIM improvement) and speech quality (0.27 UTMOS gain). These results validate the effectiveness of our speech-infilling-based paradigm and its tighter integration with the unified architecture.
For completeness, we performed ablation studies that examined the impact of different configurations of attention masks and task weighting coefficients ($\lambda$), with detailed results and analysis provided in Appendix B.
Table 2: Comparison of the zero-shot TTS capability of different TTS model variants of UniVoice.
[…]
UniVoice-TTS-speaker 5.72 0.29 3.65
UniVoice-TTS-infilling 4.66 0.56 3.92
UniVoice 4.06 0.56 3.72
Limitation
UniVoice is the first attempt in the audio field to integrate autoregression with flow matching on LLMs. It has some limitations: 1) Although it can perform speech understanding and generation tasks, it is currently limited to ASR and TTS tasks. In the future, we will try to integrate more tasks; 2) The current system is trained based on a relatively small dataset with a small language model, and there is still room for improvement in performance. We believe that using larger datasets and larger models will have better results; 3) The current system utilizes LLM to unify ASR and TTS tasks for speech understanding and generation tasks, respectively. However, the original conversational ability of LLM has not been effectively utilized, and we will expand to conversational systems in the future.
Conclusion
This paper presents UniVoice, a unified transformer framework that effectively integrates autoregressive speech recognition with flow-matching-based speech synthesis. We propose a dual attention mechanism that adaptively switches between causal and bidirectional attention patterns for ASR and TTS, respectively. Furthermore, a text-prefix-guided speech-infilling approach for high-fidelity zero-shot voice cloning is developed. The proposed unified architecture demonstrates robust performance in both speech understanding and generation tasks, establishing the viability of joint modeling through complementary paradigms. This work establishes new possibilities for end-to-end unified speech processing. Demo samples can be found at https://univoice-demo.github.io/UniVoice. To support research reproducibility in this domain, we will open-source the codes and checkpoints.
Appendix A A. Baseline Details
TTS Models
  • VALL-E (Wang et al. 2023). It use an autoregressive and an additional non-autoregressive model for discrete speech codec token generation.
  • NaturalSpeech2 (Shen et al. 2023). It use a non autoregressive model for continuous vectors generation.
  • F5-TTS [^2] https://huggingface.co/SWivid/F5-TTS (Chen et al. 2024c). It is a fully non-autoregressive text-to-speech system based on flow matching with Diffusion Transformer, which is trained on Emilia (He et al. 2024) dataset with around 100K hours Chinese and English speech dataset. We use the official checkpoint of F5-TTS for evaluation.
  • CosyVoice [^3] https://github.com/FunAudioLLM/CosyVoice (Du et al. 2024). It is a two-stage large-scale TTS system, first for autoregressive text-to-token generation, then a flow matching diffusion model for Mel-spectrogram generation. The model is trained on 170K hours of multilingual speech data. We use the official checkpoint of CosyVoice for evaluation.
  • FireRedTTS [^4] https://github.com/FireRedTeam/FireRedTTS (Guo et al. 2024a). It is a foundation TTS framework for industry-level generative speech applications. It includes an autoregressive text-to-semantic token model and a token-to waveform generation model. The system is trained with 248K hours of labeled speech data. We use the official pre-trained checkpoint to evaluate.
  • MaskGCT [^5] https://huggingface.co/amphion/MaskGCT (Wang et al. 2024b). It is large-scale non autoregressive TTS model without precise alignment information between text and speech following the mask-and-predict learning paradigm. It is trained on Emilia (He et al. 2024) dataset with around 100K hours Chinese and English speech dataset.
ASR Models
  • Whisper (Radford et al. 2023). It is an advanced ASR model developed by OpenAI, designed to transcribe speech with high accuracy. It is trained on a massive dataset of 680,000 hours of supervised audio data, covering a wide range of languages and acoustic conditions. We use Whisper-small [^6] https://huggingface.co/openai/whisper-small, Whisper-large-v2777https://huggingface.co/openai/whisper-large-v2 and Whisper-large-v3888https://huggingface.co/openai/whisper-large-v3 for evaluation.
  • Paraformer (Gao et al. 2022). Paraformer is a non-autoregressive end-to-end ASR model that achieves fast parallel decoding via CIF-based acoustic modeling. It is trained on 20K hours of English data.
  • Zipformer (Yao et al. 2023). It is a faster, more memory efficient, and better-performing transformer model for speech recognition. It is trained on 960 hours LibriSpeech dataset.
Unified Models
  • SpeechT5 (Ao et al. 2021). It is a universal pre-trained encoder-decoder model for various speech processing tasks. After pretraining, it should be fine-tuned via the loss of the downstream tasks.
  • LauraGPT (Du et al. 2023). It is a novel unified GPT-based audio LLM for audio recognition, understanding, and generation by using continuous representations for audio input and generating output audio from audio codec tokens.
  • OpusLM (Tian et al. 2025). It is a family of scalable decoder-only transformers (135M–7B parameters) that unify speech-text processing through multistream discrete token generation for both speech and text modalities. We compare the 360M and 7B models with our UniVoice.
Table 3: Comparison of performance using different $\lambda$. WER-c and WER-o represent the ASR WER evaluated on LibriSpeech test-clean dataset and test-other dataset, respectively.
[…]
$\lambda=0.01$ 4.66 0.54 3.69 4.21 7.82
$\lambda=0.005$ 4.06 0.56 3.72 3.01 6.36
Table 4: Comparison of performance using different attention masks for TTS.
[…]
AR Mask 9.85 0.49 2.23
Full Mask 4.66 0.56 3.92
Appendix B B. Experimental Result Supplements
Ablation Study on Attention Mask Strategies
We evaluated different attention mask configurations for our text-prefix-guided speech-infilling TTS approach. Table 4 shows that the use of full bidirectional attention masks consistently outperforms autoregressive masking in all evaluation metrics (WER, SIM and UTMOS). This result validates our design choice to use complete context access for high-quality speech synthesis.
Ablation Study on Different $\lambda$
We conducted experiments using different $\lambda$ to balance the ASR and TTS objectives to train our proposed UniVoice. As shown in Table 3, the set $\lambda$ to 0.005 is better than 0.01. We believe that training TTS based on flow matching is more difficult under the same transformer backbone, so it has a higher weight. However, ASR tasks are relatively simple and set with smaller weights, resulting in a greater contribution to the loss gradient of TTS during model training. This will make the model prioritize learning more difficult TTS tasks.